# assumeAssumption

### v1

@@ 4.1 Flattening trajectories
A curve is a function $x:I\to\mathbb{R}$ where $I$ is an interval of $\mathbb{R}$. We refer to trajectories as curves that solve an ODE. We say that a curve $x:I\to\mathbb{R}$ is flattening, or flattens over time, if $x(t)\prec x(s)$ for all $s,t\in I$ such that $s<t$. A curve $x:I\to\mathbb{R}$ sharpens over time if $I\ni t\mapsto x(-t)\in\mathbb{R}^{n}$ flattens over time. We seek to construct such curves using the following assumption.
Assumption 2.
@@
$\exists\omega>0:~\forall x\in U,~\forall v\in\overline{\partial}f(x),~~~\langle\nabla c(x),v\rangle=0~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}.$
The first-order condition in ˜2 induces conservation laws.
Proposition 5.
@@
    If $x:I\to U$ is a solution to $\dot{x}\in-\overline{\partial}f(x)$ almost everywhere, then $c$ is conserved.
@@
    If $x:I\to U$ is a solution to $\dot{x}=\nabla c(x)$, then $f$ is conserved.
@@
(ii) Since $f$ is definable and Lipschitz continuous, by [9, Corollary 2, Proposition 2] $f$ is path differentiable. Thus, for almost every $t\in I$,
@@
Under ˜2, gradient trajectories of the conserved quantity $c$ sharpen over time. This is because the level sets are contracting. To see why, we begin with a simple lemma.
Lemma 3.
@@
Consider the function $g:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ defined by $g(x,y)=\langle\nabla c(x)-\nabla c(y),x-y\rangle$. Since $c$ is $C^{4}$ on $U$, $g$ is $C^{3}$ on $U$. By [34, Lemma 1.2.4], there exists $M>0$ such that for all $x,y\in U$, $u\in U+\{-x\}$, and $v\in U+\{-y\}$ we have
@@
Theorem 1.
Under ˜2, for all $x_{0}\in U$ near which $f$ is nonconstant, there exists a local solution $x:[0,\overline{t})\to\mathbb{R}^{n}$ to
@@
Since $\nabla c$ is Lipschitz continuous on $U$, by [27, Theorem D.1], there exist $\overline{t},\overline{r}>0$ such that the ODE
@@
admits a unique $C^{1}$ solution $z:[0,\overline{t})\to U$ for all $z_{0}\in B_{\overline{r}}(x_{0})$. Since $\nabla c$ is $C^{2}$, by [27, Theorem D.1], the semiflow $\theta:[0,\overline{t})\times B_{\overline{r}}(x_{0})\to\mathbb{R}^{n}$ defined by $\theta(t,z_{0})=z(t)$ is $C^{2}$. By [34, Lemma 1.2.4], there exists $C>0$ such that for all $t,\widetilde{t}\in[0,\overline{t})$ and $z_{0},\widetilde{z}_{0}\in B_{\overline{r}}(x_{0})$, we have
@@
Since $f$ is locally Lipschitz definable and nonconstant near $x_{0}$, by Lemma˜1 there exist definable curves $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that $\gamma$ is $C^{1}$ on $0,\overline{r}]$ and
$\forall r\in[0,\overline{r}],~~~\gamma(r)=x_{0}+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|~~~\text{and}~~~v(r)\in\overline{\partial}f(\gamma(r)),$
@@
Let $M>0$ be given by Lemma˜3. For all $(t,r)\in[0,\overline{t})\times[0,\overline{r})$, let $x_{r}(t)=\theta(t,\gamma(r))$ and $x(t)=x_{0}(t)=\theta(t,\overline{x})$. As $\gamma$ is continuous, so is the composition $0,\overline{t})\times[0,\overline{r})\ni(t,r)\to x_{r}(t)\in\mathbb{R}^{n}$. In particular, $|x_{r}(t)-x_{0}|\leq\omega/(8\max\{L,M\})$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$ after possibly reducing $\overline{t}$ and $\overline{r}$. It follows that $|x_{r}(t)-\gamma(r)|\leq|x_{r}(t)-x_{0}|+|x_{0}-\gamma(r)|\leq\omega/(8L)+\omega/(8L)=\omega/(4L)$ and $|x_{r}(t)-x(t)|\leq|x_{r}(t)-x_{0}|+|x_{0}-x_{0}(t)|\leq\omega/(4M)$.
The main idea of the proof is now as follows. Using Proposition˜5, we have
@@
where the last inequality is due to $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|=e^{-\omega t/2}r$. The step we need to justify is the contraction. To that avail, let $y_{r}:[0,\overline{t})\to\mathbb{R}^{n}$ be defined by $y_{r}(t)=x_{r}(t)-x(t)$. Initially $y_{r}(0)=x_{r}(0)-x(0)=\gamma(r)-x_{0}=\lambda(r)v(r)\neq 0$ for all $r\in(0,\overline{r})$. Since $v(r)\in\overline{\partial}f(\gamma(r))$, we have $\langle\nabla^{2}c(\gamma(r))v(r),v(r)\rangle\leq-\omega|v(r)|^{2}$. Together with $\lambda(r)\neq 0$, we find $\langle\nabla^{2}c(\gamma(r))\lambda(r)v(r),\lambda(r)v(r)\rangle\leq-\omega|\lambda(r)v(r)|^{2}$. In other words,
@@
Each entry of $(\gamma-x_{0})/|\gamma-x_{0}|$ is definable and bounded. By the monotonicity theorem [40, 4.1], it is monotone near 0 and thus convergent. Thus there exists $u\in S^{n-1}$ such that $(\gamma(r)-x_{0})/|\gamma(r)-x_{0}|\to u$ as $r\searrow 0$. Since $\theta$ is $C^{1}$, it follows that
@@
The ODE comparison theorem [27, Theorem D.2] implies that $|y_{r}(t)|^{2}\leq e^{-\omega t}|y_{r}(0)|^{2}$, namely, $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$. As for the equivalence, it is obtained via Lemma˜1 and the change of variables $\ell=\accentset{\circ}{f}(x_{0},r)$: $\accentset{\circ}{f}(x_{0},r)\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$ iff $\overline{f}(x(t),\accentset{\circ}{f}(x_{0},r))\leq\overline{f}(x(t),\accentset{\circ}{f}(x(t),e^{-\omega t/2}r))$ iff $\overline{f}(x(t),\ell)\leq e^{-\omega t/2}\overline{f}(x_{0},\ell)$. ∎
Due to the local nature of Theorem˜1, it does not guarantee that reversing time yields flattening trajectories from a given initial point. In order to do so, we will add some assumptions and rely on the following fact.
Fact 14.
@@
Corollary 5.
Under ˜2, let $x:I\to U$ be such that $\dot{x}=-\nabla c(x)$ on a compact interval $I$ of $\mathbb{R}$ containing $t_{0}$. Let $s,t\in I$ be such that $s<t$.
@@
(i) Let $t\in I$. By Theorem˜1 and Proposition˜3, $\mathrm{lip}f(x(t))+o(1)=\accentset{\circ}{f}(x(t),r)/r\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r=e^{-\omega(t-s)/2}\mathrm{lip}f(x(s))+o(1)$ for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$. ˜14 enables one to extend it to any $s\in I\cap(-\infty,t]$. (ii) Assuming the existence of a solution $x:I\to U$ to the ODE passing through $x_{0}=x(t_{0})$ at time $t_{0}$ implies the existence of solutions on an open interval $J\supset I$ for initial points in a neighborhood $U_{0}$ of $x_{0}$. This holds because $\nabla c$ is Lipschitz continuous on $U$, which also implies uniqueness of solutions (using the ODE comparison theorem [27, Theorem D.2]). Thus the flow $\theta:J\times U_{0}\to U$ is well defined and $C^{k}$. For any $t\in J$, $\theta_{t}=\theta(t,\cdot)$ defines a diffeomorphism from $U_{0}$ to $\theta_{t}(U_{0})$ [27, p. 209], which is a neighborhood of $x(t)$. Conservation implies that $f=f\circ\theta_{t}$. By the chain rule, $(f\circ\theta_{t})^{\prime}(x)=f^{\prime}(\theta_{t}(x))\circ\theta_{t}^{\prime}(x)$. In particular, $0=f^{\prime}(x_{0})(v)=(f\circ\theta_{t})^{\prime}(x_{0})(v)=f^{\prime}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)$ for all $v\in\mathbb{R}^{n}$, so that $f^{\prime}(\theta_{t}(x_{0}))=0$. By induction, $(f\circ\theta_{t})^{(i)}(x_{0})v^{i}=f^{(i)}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)^{i}$ for all $i\in\llbracket 1,k\rrbracket$ and thus $f^{(i)}(x(t))=0$ for all $i\in\llbracket 1,k-1\rrbracket$. It follows that $\|f^{(k)}(x(t))\|/k!+o(1)=\accentset{\circ}{f}(x(t),r)/r^{k}\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r^{k}=e^{-k\omega(t-s)/2}\|f^{(k)}(x(s))\|/k!+o(1)$ for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$ by Theorem˜1 and Proposition˜4. One again concludes by ˜14. ∎
Remark 2.
@@
One can actually weaken the equality $\langle\nabla c(x),v\rangle=0$ for all $v\in\overline{\partial}f(x)$ in ˜2 to an inequality $\langle\nabla c(x),v\rangle\geq 0$. In exchange, one must assume that $x(\cdot)$ is defined on $0,\overline{t}]$ (at no cost) and $x(\overline{t})\in\arg\min\{f(x):x\in U\}$ in Theorem˜1. In the proof, $f$ instead increases along the trajectories of $\nabla c$, but is still constant on $x(\cdot)$, so that $|f(x_{r}(0))-f(x(0))|\leq|f(x_{r}(t))-f(x(t))|$.
@@
Similarly, one must assume that $x(t_{0})\in\arg\min\{f(x):x\in U\}$ in Corollary˜5. In the proof, instead of the chain rule, one uses Taylor expansions: $f(y)=f(x_{0})+\langle\nabla f(x_{0}),y-x_{0}\rangle+o(|y-x_{0}|)$ and $f(\theta_{t}(y))=f(\theta_{t}(x_{0}))+\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)$. Since $f(x_{0})=f(\theta_{t}(x_{0}))\leq f(\theta_{t}(y))\leq f(y)$, this yields $\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)=o(|y-x_{0}|)$ and so $\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}^{\prime}(x_{0})v\rangle=0$ for all $v\in\mathbb{R}^{n}$, namely $\nabla f(x(t))=0$. The rest follows by induction.
@@
With these new assumptions, it is sufficient to satisfy the second inequality in ˜2, i.e., $\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}$ for all $v\in\overline{\partial}f(x)$ and $x\in U$, merely for all $x\in U\setminus\arg\min\{f(x):x\in U\}$.

### v2

@@ 4.1 Flattening trajectories
A curve is a function $x:I\to\mathbb{R}$ where $I$ is an interval of $\mathbb{R}$.
Definition 4.1.
A trajectory of $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is an absolutely continuous curve $x:I\to\mathbb{R}^{n}$ where $I$ is an interval of $\mathbb{R}$ such that
[…]
If $F$ is singleton-valued and continuous, then its trajectories are $C^{1}$. We refer to solutions to the differential inclusion $\dot{x}\in F(x)$ as trajectories of $F$ over an interval $I=[0,T)$ for some $T\in(0,\infty]$ or $I=[0,T]$ for some $T\in(0,\infty)$. A solution $x:I\to\mathbb{R}^{n}$ is maximal if for any other solution $y:J\to\mathbb{R}^{n}$ such that $I\subseteq J$ and $x(t)=y(t)$ for all $t\in I$, we have $I=J$. A solution is globally defined if $I=[0,\infty)$.
Definition 4.2.
Given $f:\mathbb{R}^{n}\to\mathbb{R}$, a curve $x:I\to\mathbb{R}^{n}$ is flattening, or flattens over time (resp. strictly), if $x(t)\preceq x(s)$ (resp. $x(t)\prec x(s)$) for all $s,t\in I$ such that $s<t$. A curve $x:I\to\mathbb{R}$ sharpens (resp. strictly) over time if $I\ni t\mapsto x(-t)\in\mathbb{R}^{n}$ flattens over time (resp. strictly).
We seek to construct such flattening curves using the following assumption. It is satisfied in Example˜5.6, [25, Examples 4.4, 4.5], and [27, Examples 4.2, 4.3, …, 4.8].
Assumption 4.3.
@@
$\exists\omega>0:~\forall x\in U,~\forall v\in\overline{\nabla}f(x),~~~\langle\nabla c(x),v\rangle=0~~\text{and}~~~\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}.$
The first-order condition in Assumption˜4.3 induces conservation laws.
Proposition 4.4.
@@
    If $x:I\to U$ is a trajectory of $\overline{\partial}f$, then $c$ is conserved.
@@
    If $x:I\to U$ is a trajectory of $\nabla c$, then $f$ is conserved.
@@
(ii) Since $f$ is definable and Lipschitz continuous, by the chain rule in [16, Corollary 5.4], for almost every $t\in I$,
@@
Under Assumption˜4.3, gradient trajectories of the conserved quantity $c$ strictly sharpen over time. This is because the level sets are contracting. To see why, we begin with a simple lemma.
Lemma 4.5.
@@
Consider the function $g:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ defined by $g(x,y)=\langle\nabla c(x)-\nabla c(y),x-y\rangle$. Since $c$ is $C^{4}$ on $U$, $g$ is $C^{3}$ on $U$. By [39, Lemma 1.2.4], there exists $M>0$ such that for all $x,y\in U$, $u\in U+\{-x\}$, and $v\in U+\{-y\}$ we have
@@
Theorem 4.6.
Under Assumption˜4.3, for all $x_{0}\in U$ near which $f$ is nonconstant, there exists a solution $x:[0,\overline{t})\to\mathbb{R}^{n}$ to
@@
Since $\nabla c$ is Lipschitz continuous on $U$, by [32, Theorem D.1], there exist $\overline{t},\overline{r}>0$ such that the ODE
@@
admits a unique $C^{1}$ solution $z:[0,\overline{t})\to U$ for all $z_{0}\in B_{\overline{r}}(x_{0})$. Since $\nabla c$ is $C^{2}$, by [32, Theorem D.1], the semiflow $\theta:[0,\overline{t})\times B_{\overline{r}}(x_{0})\to\mathbb{R}^{n}$ defined by $\theta(t,z_{0})=z(t)$ is $C^{2}$. By [39, Lemma 1.2.4], there exists $C>0$ such that for all $t,\widetilde{t}\in[0,\overline{t})$ and $z_{0},\widetilde{z}_{0}\in B_{\overline{r}}(x_{0})$, we have
@@
Since $f$ is locally Lipschitz, regular, definable, and nonconstant near $x_{0}$, by Lemma˜3.13 there exist definable curves $\gamma:[0,\overline{r}]\to\mathbb{R}^{n}$, $\lambda:[0,\overline{r}]\to\mathbb{R}$, and $v:[0,\overline{r}]\to\mathbb{R}^{n}$ such that $\gamma$ is $C^{1}$ on $0,\overline{r}]$ and
$\forall r\in[0,\overline{r}],~~\gamma(r)=x_{0}+\lambda(r)v(r)\in\arg\max_{\overline{B}_{r}(x_{0})}|f-f(x_{0})|,~~|\gamma(r)-x_{0}|=r,~~v(r)\in\overline{\nabla}f(\gamma(r)),$
@@
Let $M>0$ be given by Lemma˜4.5. For all $(t,r)\in[0,\overline{t})\times[0,\overline{r})$, let $x_{r}(t)=\theta(t,\gamma(r))$ and $x(t)=x_{0}(t)=\theta(t,\overline{x})$. As $\gamma$ is continuous, so is the composition $0,\overline{t})\times[0,\overline{r})\ni(t,r)\to x_{r}(t)\in\mathbb{R}^{n}$. In particular, $|x_{r}(t)-x_{0}|\leq\omega/(8\max\{L,M\})$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$ after possibly reducing $\overline{t}$ and $\overline{r}$. It follows that $|x_{r}(t)-\gamma(r)|\leq|x_{r}(t)-x_{0}|+|x_{0}-\gamma(r)|\leq\omega/(8L)+\omega/(8L)=\omega/(4L)$ and $|x_{r}(t)-x(t)|\leq|x_{r}(t)-x_{0}|+|x_{0}-x_{0}(t)|\leq\omega/(4M)$.
The main idea of the proof is now as follows. Using Proposition˜4.4, we have
@@
where the last inequality is due to $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|=e^{-\omega t/2}r$. The step we need to justify is the contraction. To that avail, let $y_{r}:[0,\overline{t})\to\mathbb{R}^{n}$ be defined by $y_{r}(t)=x_{r}(t)-x(t)$. Initially $y_{r}(0)=x_{r}(0)-x(0)=\gamma(r)-x_{0}=\lambda(r)v(r)\neq 0$ for all $r\in(0,\overline{r})$. Since $v(r)\in\overline{\nabla}f(\gamma(r))$, by Assumption˜4.3 we have $\langle\nabla^{2}c(\gamma(r))v(r),v(r)\rangle\leq-\omega|v(r)|^{2}$. Together with $\lambda(r)\neq 0$, we find $\langle\nabla^{2}c(\gamma(r))\lambda(r)v(r),\lambda(r)v(r)\rangle\leq-\omega|\lambda(r)v(r)|^{2}$. In other words,
@@
Each entry of $(\gamma-x_{0})/|\gamma-x_{0}|$ is definable and bounded. By the monotonicity theorem [48, 4.1], it is monotone near 0 and thus convergent. Thus there exists $u\in S^{n-1}$ such that $(\gamma(r)-x_{0})/|\gamma(r)-x_{0}|\to u$ as $r\searrow 0$. Since $\theta$ is $C^{1}$, it follows that
@@
The ODE comparison theorem [32, Theorem D.2] implies that $|y_{r}(t)|^{2}\leq e^{-\omega t}|y_{r}(0)|^{2}$, namely, $|x_{r}(t)-x(t)|\leq e^{-\omega t/2}|x_{r}(0)-x(0)|$ for all $t\in[0,\overline{t})$ and $r\in[0,\overline{r})$. As for the equivalence, it is obtained via Lemma˜3.13 and the change of variables $\ell=\accentset{\circ}{f}(x_{0},r)$: $\accentset{\circ}{f}(x_{0},r)\leq\accentset{\circ}{f}(x(t),e^{-\omega t/2}r)$ iff $\overline{f}(x(t),\accentset{\circ}{f}(x_{0},r))\leq\overline{f}(x(t),\accentset{\circ}{f}(x(t),e^{-\omega t/2}r))$ iff $\overline{f}(x(t),\ell)\leq e^{-\omega t/2}\overline{f}(x_{0},\ell)$. ∎
Due to the local nature of Theorem˜4.6, it does not guarantee that reversing time yields strictly flattening trajectories from a given initial point. In order to do so, we will add some assumptions and rely on the following fact.
Fact 4.7.
@@
Corollary 4.8.
Under Assumption˜4.3, let $x:I\to U$ be such that $\dot{x}=-\nabla c(x)$ on a compact interval $I$ of $\mathbb{R}$ containing $t_{0}$. Let $s,t\in I$ be such that $s<t$.
@@
(i) Let $t\in I$. By Theorem˜4.6 and Proposition˜3.15,
$\mathrm{lip}f(x(t))+o(1)=\accentset{\circ}{f}(x(t),r)/r\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r=e^{-\omega(t-s)/2}\mathrm{lip}f(x(s))+o(1)$
for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$. Fact˜4.7 enables one to extend it to any $s\in I\cap(-\infty,t]$.
@@
(ii) Assuming the existence of a solution $x:I\to U$ to the ODE passing through $x_{0}=x(t_{0})$ at time $t_{0}$ implies the existence of solutions on an open interval $J\supseteq I$ for initial points in a neighborhood $U_{0}$ of $x_{0}$. This holds because $\nabla c$ is Lipschitz continuous on $U$, which also implies uniqueness of solutions (using the ODE comparison theorem [32, Theorem D.2]). Thus the flow $\theta:J\times U_{0}\to U$ is well defined and $C^{k}$. For any $t\in J$, $\theta_{t}=\theta(t,\cdot)$ defines a diffeomorphism from $U_{0}$ to $\theta_{t}(U_{0})$ [32, p. 209], which is a neighborhood of $x(t)$. Conservation implies that $f=f\circ\theta_{t}$. By the chain rule, $(f\circ\theta_{t})^{\prime}(x)=f^{\prime}(\theta_{t}(x))\circ\theta_{t}^{\prime}(x)$. In particular,
$\forall v\in R^{n},~~~0=f^{\prime}(x_{0})(v)=(f\circ\theta_{t})^{\prime}(x_{0})(v)=f^{\prime}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v),$
so that $f^{\prime}(\theta_{t}(x_{0}))=0$. By induction, $(f\circ\theta_{t})^{(i)}(x_{0})v^{i}=f^{(i)}(\theta_{t}(x_{0}))(\theta_{t}^{\prime}(x_{0})v)^{i}$ for all $i\in\llbracket 1,k\rrbracket$ and thus $f^{(i)}(x(t))=0$ for all $i\in\llbracket 1,k-1\rrbracket$. It follows that
$\|f^{(k)}(x(t))\|/k!+o(1)$ $=\accentset{\circ}{f}(x(t),r)/r^{k}\leq\accentset{\circ}{f}(x(s),e^{-\omega(t-s)/2}r)/r^{k}$
$=e^{-k\omega(t-s)/2}\|f^{(k)}(x(s))\|/k!+o(1)$
for all $s\in I\cap(-\infty,t]$ sufficiently close to $t$ by Theorem˜4.6 and Proposition˜3.17. One again concludes by Fact˜4.7. ∎
Remark 4.9.
@@
One can actually weaken the equality $\langle\nabla c(x),v\rangle=0$ for all $v\in\overline{\nabla}f(x)$ in Assumption˜4.3 to an inequality $\langle\nabla c(x),v\rangle\geq 0$. In exchange, one must assume that $x(\cdot)$ is defined on $0,\overline{t}]$ (at no cost) and $x(\overline{t})\in\arg\min\{f(x):x\in U\}$ in Theorem˜4.6. In the proof, $f$ instead increases along the trajectories of $\nabla c$, but is still constant on $x(\cdot)$, so that $|f(x_{r}(0))-f(x(0))|\leq|f(x_{r}(t))-f(x(t))|$.
@@
Similarly, one must assume that $x(t_{0})\in\arg\min\{f(x):x\in U\}$ in Corollary˜4.8. In the proof, instead of the chain rule, one uses Taylor expansions: $f(y)=f(x_{0})+\langle\nabla f(x_{0}),y-x_{0}\rangle+o(|y-x_{0}|)$ and $f(\theta_{t}(y))=f(\theta_{t}(x_{0}))+\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)$. Since $f(x_{0})=f(\theta_{t}(x_{0}))\leq f(\theta_{t}(y))\leq f(y)$, this yields $\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}(y)-\theta_{t}(x_{0})\rangle+o(|\theta_{t}(y)-\theta_{t}(x_{0})|)=o(|y-x_{0}|)$ and so $\langle\nabla f(\theta_{t}(x_{0})),\theta_{t}^{\prime}(x_{0})v\rangle=0$ for all $v\in\mathbb{R}^{n}$, namely $\nabla f(x(t))=0$. The rest follows by induction.
@@
With these new assumptions, it is sufficient to satisfy the second inequality in Assumption˜4.3, i.e., $\langle\nabla^{2}c(x)v,v\rangle\leq-\omega|v|^{2}$ for all $v\in\overline{\nabla}f(x)$ and $x\in U$, merely for all $x\in U\setminus\arg\min\{f(x):x\in U\}$.